# Rethinking Graph Structure Learning in the Era of LLMs

### v4

@@ Abstract
Recently, the emergence of LLMs has prompted researchers to integrate language descriptions into graphs, aiming to enhance model encoding capabilities from a data-centric perspective. This graph representation is called text-attributed graphs (TAGs). A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm. Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLM? (2) How can we design an efficient model architecture that enables seamless integration of LLM for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 11 datasets demonstrate that LLaTA enjoys flexibility-incorporated with any backbone; scalability-outperforms other LLM-enhanced graph learning methods; effectiveness-achieves SOTA predictive performance.
@@ 1 Introduction
In recent years, the rise of LLMs in graph ML [30, 3] has led to the emergence of text-attributed graphs (TAGs) as a novel data representation. This data structure leverages textual information to provide fine-grained descriptions of graphs, driving a significant shift in data-centric graph learning [14]. Inspired by this, efficiently and robustly learning TAGs has become an urgent priority.
To achieve this, graph structure learning (GSL) is a promising approach to enhancing data utility for vanilla graph neural networks (GNNs). However, most existing GSL methods [39] are designed for traditional graphs and cannot effectively process rich textual information, leading to suboptimal performance. Despite recent efforts in LLM-based GSL approaches, such as GraphEdit [7] and LLM4RGNN [34], their optimization objectives and model architectures still follow traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL methods [36, 29, 28, 34] rely on carefully designed loss functions to maintain a graph learner (i.e., structure optimizer) coupled with well-defined instruction datasets or specific GNN backbones (i.e., improved structure tailored for the downstream encoder). Based on this, we make two key observations and aim to explore them to guide the development of a new GSL paradigm in the era of LLMs.
Observation 1 (Optimization Objective): The graph learner often serves as an edge predictor, heavily relying on end-to-end training with a specific downstream backbone. In the era of LLMs, LLMs will play a crucial role in graph learners. In this context, performing full-parameter LLM training to obtain an edge predictor is infeasible. While two recent fine-tuning methods have been proposed, the complexity of constructing instruction datasets and tuning still hinders their efficiency. This leads to Question 1: How can we define a reasonable optimization objective for LLM-enhanced GSL?
@@
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which seamlessly integrates both topology and text insights by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure, with complete theoretical support. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based methods by 1.3%-2.5% in accuracy while running 2.5h-9.2h faster.
@@ 2.1 Notations and Problem Formulation
[…]
@@ 2.2 Complexity Metrics of Graph Topology
Structural Entropy. Motivated by Shannon entropy [22], structural entropy (SE) [12] is an effective measurement for quantifying the dynamic complexity of graph topology. By minimizing SE, we can reduce structure uncertainty and capture inherent patterns, thereby supporting downstream tasks in a robust and interpretable manner. In other words, this approach uncovers meaningful structural patterns and ensures that it aligns more effectively with task-specific requirements. This has made it a pivotal tool for GNNs, gaining significant attention in recent years.
Structural Encoding Tree. By minimizing SE using the greedy algorithm, we can construct a structural encoding tree $\mathcal{\mathcal{T}}$. This tree reorganizes the original graph and simulates the natural evolution of the graph through its inherent hierarchical structure. Specifically, its leaf nodes correspond to the original graph nodes, and non-leaf nodes represent multi-level communities. Low-level communities capture connected leaf nodes with high homophily, while higher levels reflect unseen structural patterns. It offers GSL a new perspective. More details can be found in Appendix A-B.
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods utilize the structural encoding tree for various graph tasks, such as node clustering [21], community detection [16], multi-layer coarsening [27], and embedding dimension estimation [31]. Recent SEGSL [42] improves edge connectivity and quality with SE-based encoding tree, but it still relies on end-to-end tree-based training without incorporating LLMs.
Traditional GSL. These methods can be categorized into metric-based [32, 33, 13], probabilistic sampling [37, 18, 15], and directly learnable methods [10, 17]. Specifically, they refine graph structure using well-designed measurement (e.g., homophily and connectivity), edge re-weight, random sampling, or direct topology optimization without textual information.
LLM-based GSL. GraphEdit [7] improves node connectivity by adding neighbors through a edge predictor and refines the structure using a fine-tuned LLM. LLM4RGNN [34] leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. In contrast, LangGSL [24] integrates LLMs with graph structure learning to jointly optimize node features and graph structure, improving performance across various tasks. In Appendix C.2, we detail the advantages of LLaTA over SEGSL and LLM-based methods.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, existing GSL methods aim to train an effective edge predictor as the graph learner. This typically requires coupling the graph learner with a specific backbone and jointly optimizing them. Although effective, this objective becomes inefficient for LLMs due to their large number of parameters. Our investigation shows that current LLM-based GSL approaches often adopt instruction fine-tuning to mitigate full-parameter training. However, understanding complex graph structures and crafting appropriate instruction datasets can be labor-intensive, as reflected in their elaborate workflows [7, 34]. To address this, we reformulate GSL and introduce a tree-based optimization framework that avoids edge predictors, as illustrated in Fig. 1 (a)-(c). The core idea is to develop a language-aware tree sampler for efficient GSL. Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for implementation of the above framework in LLaTA.
@@ 3.2 Model Architecture Review
As outlined by Obs. 2 in Sec. 1, integrating LLMs in GSL remains a challenge. To establish model design principles for tree optimization, we revisit existing GSL paradigms. The Coupled Paradigm tightly integrates the graph learner and backbone for task-specific learning. However, it faces limitations such as unstable performance due to restricted generalizability and module dependent, and inflexible deployment as switching tasks or backbones requires retraining. In contrast, the Decoupled Paradigm trains the graph learner and backbone independently, offering better compatibility with LLMs for GSL. Although recent LLM-based GSL methods adopt this approach, they still face challenges such as indirect integration of topology and text, reliance on instruction datasets, and the complexity of fine-tuning. For a more efficient and reliable approach, we recommend a fully decoupled paradigm, reducing fine-tuning reliance and focusing on high-quality in-context prompts. Detailed discussion about the GSL paradigms is provided in the Appendix C.1.
To achieve this, we leverage the in-context learning ability of LLMs to jointly understand graph topology and node text for reliable inference. Specifically, (a) Tree Construction: we first build a structural encoding tree to generate topology-aware in-context prompts, enhancing the LLM’s understanding of graph structure. Then, (b) Tree Optimization: based on these prompts, we use LLMs to infer semantic relationships among nodes from their textual descriptions, and reassign leaf dependencies to optimize the tree structure based on the community structure. The resulting hierarchical tree integrates both structural and textual information, forming a strong basis for GSL. Finally, (c) Improved Structure: we apply a leaf-oriented two-step sampling to select the current node and a candidate set for edge modification, yielding an improved graph for any downstream scenario. For deeper insights into the leaf-oriented sampler within GSL, please refer to Appendix B.
@@ 4 Our Method
We instantiate the three components outlined in Fig. 1 and propose LLaTA, shown in Fig. 2. The correspondences are: (a)-(1) Topology-aware In-context, (b)-(2) Tree-prompted LLM Inference and (c)-(3) Leaf-oriented Two-step Sampling.
@@ 4.1 Topology-aware In-context Construction
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by [12, 42]. The optimization objective definition:
$\mathcal{T}^{\star}\!\!=\!\arg\!\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(\mathcal{G}),\>\>\>\>\mathcal{H}^{\mathcal{T}}(\mathcal{G})=\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\!\mathcal{H}^{\mathcal{T}}(\mathcal{G},\phi)=-\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\frac{g_{\phi}}{\operatorname{vol}(\mathcal{G})}\log\!\frac{\operatorname{vol}\left(\phi\right)}{\operatorname{vol}\left(\phi^{+}\right)},$ (1)
@@
Theorem 2 (Implicit Global Constraints in Low-Level Communities).
@@
Drawing on Theorems 1 and  2, we infer that low-level communities in the encoding tree provide a more faithful representation of the graph’s topology, as they simultaneously preserve local connectivity and global structural constraints. Accordingly, for each leaf node $\alpha$, we assign its associated low-level community $\mathcal{C}^{\ell}_{\alpha^{+}}$ as a topological context to enhance the accuracy of downstream inference by the LLMs. The proof process of Theorems 1 and  2 is provided in Appendix F.1-F.2.
@@ 4.2 Tree-prompted LLM Inference
Motivation: Based on the structural encoding tree $\mathcal{T}$, we further integrate node language descriptions to enable comprehensive in-context learning with LLMs, fostering a better understanding of both topology and textual information in the original graph. This allows LLMs to make reliable inferences without fine-tuning. These inferences are then used to optimize the encoding tree, enhancing homophily within low-level communities and laying a strong foundation for subsequent GSL.
@@
$t^{\star}_{\alpha}=\operatorname{Concat}\left(\left\{t_{\alpha}\}\cup\{t_{j}\mid j\in\text{Top}_{k}\left(\left\{w_{\alpha\beta}\mid\beta\in\mathcal{C}^{\ell}_{\alpha^{+}}\right\}\right),w_{\alpha\beta}\geq\epsilon\right\}\right),$ (2)
$w_{\alpha\beta}=\frac{\exp\left(\operatorname{sim}(x_{\alpha},x_{\beta})\right)}{\sum\limits_{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}\exp\left(\operatorname{sim}(x_{\alpha},x_{\gamma})\right)},\>\>\>\>\operatorname{sim}(x_{\alpha},x_{\beta})=\frac{\sum_{i=1}^{f\text{-dim}}x_{\alpha}^{i}\cdot x_{\beta}^{i}}{\sqrt{\sum_{i=1}^{f\text{-dim}}(x_{\alpha}^{i})^{2}\sum_{i=1}^{f\text{-dim}}(x_{\beta}^{i})^{2}}}.$ (3)
Community of Thought. Building on the enhanced leaf nodes, we propose a Community of Thought (CoT) prompt mechanism, where each node is represented together with its community-aware descriptions and related neighbors. This design enables LLMs to jointly capture topological and semantic information, thereby supporting more reliable inference. Through this mechanism, the LLM can (i) infer the probability distribution over label classes for each node, and (ii) generate high-quality embeddings within its semantic space. These embeddings are then employed for fine-grained optimization of the structural encoding tree. In summary, CoT facilitates efficient in-context learning for LLMs by leveraging the structure-aware encoding tree. Formally, given a target leaf node $\alpha$, the LLM inference with CoT can be expressed as:
$\mathbf{z}_{\alpha}=\Psi\!\left(\mathrm{LLM\mbox{-}CoT}\!\left(\mathcal{D}_{\text{task}},\mathcal{D}_{\text{topo}},\mathcal{D}_{\text{sema}},t^{\star}_{\alpha}\right)\right),\quad\Psi:\text{ANS}_{\alpha}\mapsto\mathbb{R}^{N_{\text{class}}},$ (4)
$y^{cls}_{\alpha}\sim\mathrm{Categorical}\!\left(p_{\theta}(y\mid\mathbf{z}_{\alpha})\right),\quad p_{\theta}(y=i\mid\mathbf{z}_{\alpha})=\frac{\exp(z_{\alpha,i})}{\sum_{j=1}^{N_{\text{class}}}\exp(z_{\alpha,j})},\;\;i=1,\ldots,N_{\text{class}},$ (5)
where $\Phi(\cdot)$ is the extraction function that maps the LLM’s textual answer $\text{ANS}_{\alpha}$ into a numerical logit vector $\mathbf{z}_{\alpha}\in\mathbb{R}^{N_{\text{class}}}$, $y^{cls}_{\alpha}\in\Delta^{N_{\text{class}}-1}$ denotes the soft label probability simplex for node $\alpha$. Here, $\mathcal{D}_{\text{task}}$ specifies the task, while $\mathcal{D}_{\text{topo}}$ and $\mathcal{D}_{\text{sema}}$ encode topological and semantic context, respectively. An illustration of prompt construction is provided in Fig. 2.
@@
Based on Theorem 3, we have that the similarity of soft labels $y^{cls}$ can approximate the similarity of ground-truth labels $y$ with controllable bias $\varepsilon^{y}(\alpha\beta)$. This provides robust theoretical guarantee and support for the subsequent Leaf Dependency Allocation and Semantic Similarity Sampling.
@@
$\mathcal{P}^{\star}=\arg\max_{\mathcal{P}}\Bigg\{\sum_{\mathcal{C}\in\mathcal{P}}\max_{c}\sum_{v\in\mathcal{C}}y^{cls}_{v,c}+\beta\,\sum_{\mathcal{C}\in\mathcal{P}}\mathrm{Sil}(\mathcal{C};\mathbf{Y}^{cls})\Bigg\},\;\;\mathcal{T}^{\star}=\mathrm{UpdateTree}(\mathcal{T},\mathcal{P}^{\star}).$ (6)

### v5

# Unlocking Graph Structure Learning with Tree-Guided Large Language Models

@@ Abstract
Recently, the emergence of large language models (LLMs) has motivated integrating language descriptions into graphs, forming text-attributed graphs (TAGs) that enhance model encoding capabilities from a data-centric perspective. A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm. Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLMs? (2) How can we design an efficient model architecture that enables seamless integration of LLMs for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 11 datasets demonstrate that LLaTA enjoys flexibility—incorporated with any backbone; scalability—outperforms other LLM-based GSL methods; and effectiveness—achieving SOTA predictive performance across a variety of datasets from different domains.
Graph Structure Learning, Large Language Model, ICML
@@ 1 Introduction
Recently, the rise of LLMs in graph ML (Yan et al., 2023; Chen et al., 2024) has led to the development of text-attributed graphs (TAGs), which leverage textual information for fine-grained graph descriptions, driving a shift in data-centric graph learning (Li et al., 2024). This has made efficiently and robustly learning TAGs an urgent priority.
To achieve this, graph structure learning (GSL) is a promising approach to enhancing data utility for vanilla graph neural networks (GNNs). However, most existing GSL methods (Zhiyao et al., 2024) are designed for traditional graphs and cannot effectively process rich textual information, leading to suboptimal performance. Despite recent efforts in LLM-based GSL approaches, such as GraphEdit (Guo et al., 2024) and LLM4RGNN (Zhang et al., 2024), their optimization objectives and model architectures still follow traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL methods (Zhao et al., 2021; Wu et al., 2022b, 2023; Zhang et al., 2024) rely on carefully designed loss functions to couple a graph learner (i.e., structure optimizer) with specific instruction datasets or GNN backbones, improving the structure for the downstream encoder. Based on this, we make two key observations and aim to explore them to guide the development of a new GSL paradigm in the era of LLMs.
Observation 1 (Optimization Objective): Graph learners typically function as edge predictors, relying heavily on end-to-end training with a downstream backbone. With the rise of LLMs, they are expected to play a key role in graph learning. However, full-parameter LLM training for edge prediction is infeasible. While recent fine-tuning methods exist, constructing instruction datasets and performing tuning remain complex and inefficient. This leads to Question 1: How can we define a suitable optimization objective for LLM-enhanced graph structure learning?
@@
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which integrates both topology and text by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure, with complete theoretical support. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based methods by 1.3%-2.5% in accuracy while running 2.5h-9.2h faster.
@@ 2.1 Notations and Problem Formulation
Graph Structure Learning in TAGs. In this paper, we aim to learn an improved graph structure $\mathbf{A}^{\star}$ that jointly exploits the original topology $\mathbf{A}$ and textual information $\mathrm{T}$ for arbitrary downstream tasks. Given predictions $\hat{\mathbf{Y}}$ from a backbone operating on $\mathbf{A}^{\star}$, we optimize $\mathcal{L}_{\text{task}}(\hat{\mathbf{Y}},\mathbf{Y})+\alpha\,\mathcal{L}_{\text{reg}}\big(\mathbf{A}^{\star};\mathbf{A},\mathrm{T}\big),$ where $\mathcal{L}_{\text{task}}$ is the task-specific loss (e.g., node classification), and $\mathcal{L}_{\text{reg}}$ regularizes the learned structure using both topology and text.
@@ 2.2 Complexity Metrics of Graph Topology
Structural Entropy. Motivated by Shannon entropy (Shannon, 1948), structural entropy (SE) (Li and Pan, 2016) provides an effective measure of the dynamic complexity of graph topology. By minimizing SE, we reduce structural uncertainty and reveal inherent patterns that support downstream tasks in a robust and interpretable way. As a result, SE has become a pivotal tool for GNNs and has attracted increasing attention in recent years.
Structural Encoding Tree. By greedily minimizing SE, we construct a structural encoding tree $\mathcal{T}$ that hierarchically reorganizes the original graph. Its leaf nodes correspond to the original graph nodes, while internal nodes represent multi-level communities: lower levels group highly homophilous nodes, and higher levels capture more global structural patterns. This hierarchical view offers a new perspective for GSL. More details can be found in Appendix AB.
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods utilize the structural encoding tree for various graph tasks, such as node clustering (Pan et al., 2021), community detection (Liu et al., 2019), multi-layer coarsening (Wu et al., 2022a), and embedding dimension estimation (Yang et al., 2023). Recent SEGSL (Zou et al., 2023) improves edge connectivity and quality with SE-based encoding tree, but it still relies on end-to-end tree-based training without incorporating LLMs.
Traditional GSL. These methods can be categorized into metric-based (Yu et al., 2020; Zhang and Zitnik, 2020; Li et al., 2022), probabilistic sampling (Zheng et al., 2020; Luo et al., 2021; Liu et al., 2022a), and directly learnable methods (Jin et al., 2020; Liu et al., 2022b). Specifically, they refine graph structure using well-designed measurement (e.g., homophily and connectivity), edge re-weight, random sampling, or direct topology optimization without textual information.
LLM-based GSL. GraphEdit (Guo et al., 2024) improves node connectivity by adding neighbors through a edge predictor and refines the structure using a fine-tuned LLM. LLM4RGNN (Zhang et al., 2024) leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. In contrast, LangGSL (Su et al., 2024) integrates LLMs with graph structure learning to jointly optimize node features and graph structure, improving performance across various tasks. In Appendix C.2, we detail the advantages of LLaTA over SEGSL and LLM-based methods.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, existing GSL methods aim to train an effective edge predictor as the graph learner. This typically requires coupling the graph learner with a specific backbone and jointly optimizing them. Although effective, this objective becomes inefficient for LLMs due to their large number of parameters. Our investigation shows that current LLM-based GSL approaches often adopt instruction fine-tuning to mitigate full-parameter training. However, understanding complex graph structures and crafting appropriate instruction datasets can be labor-intensive, as reflected in their elaborate workflows (Guo et al., 2024; Zhang et al., 2024). To address this, we reformulate GSL and introduce a tree-based optimization framework that avoids edge predictors, as illustrated in Fig. 1 (a)-(c). The core idea is to develop a language-aware tree sampler for efficient GSL. Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for implementation of the above framework in LLaTA.
@@ 3.2 Model Architecture Review
As outlined by Obs. 2 in Sec. 1, integrating LLMs in GSL remains a challenge. To establish model design principles for tree optimization, we revisit existing GSL paradigms. The Coupled Paradigm tightly integrates the graph learner and backbone for task-specific learning. However, it faces limitations such as unstable performance due to restricted generalizability and module dependent, and inflexible deployment as switching tasks or backbones requires retraining. In contrast, the Decoupled Paradigm trains the graph learner and backbone independently, offering better compatibility with LLMs for GSL. Although recent LLM-based GSL methods adopt this approach, they still face challenges such as indirect integration of topology and text, reliance on instruction datasets, and the complexity of fine-tuning. For a more efficient and reliable approach, we recommend a fully decoupled paradigm, reducing fine-tuning reliance and focusing on high-quality in-context prompts. A detailed discussion of GSL paradigms is given in Appendix C.1.
To achieve this, we leverage the in-context learning ability of LLMs to jointly understand graph topology and node text for reliable inference. Specifically, (a) Tree Construction: we first build a structural encoding tree to generate topology-aware in-context prompts, enhancing the LLM’s understanding of graph structure. Then, (b) Tree Optimization: given these prompts, the LLM infers semantic relationships among nodes from their textual descriptions and reassigns leaf dependencies to refine the community-based tree. The resulting hierarchical tree jointly encodes structural and textual information, providing a strong basis for GSL. Finally, (c) Improved Structure: we apply a leaf-oriented two-step sampling to select the current node and a candidate set for edge modification, yielding an improved graph for any downstream scenario. For deeper insights into the leaf-oriented sampler in GSL, see Appendix B.
@@ 4 Our Method
We instantiate the three conceptual components outlined in Fig. 1 into a concrete framework, LLaTA, whose overall pipeline is illustrated in Fig. 2. In this framework, (a) corresponds to (1) topology-aware in-context construction, (b) to (2) tree-prompted LLM inference, and (c) to (3) leaf-oriented two-step sampling, which together form the complete structure-learning pipeline.
@@ 4.1 Topology-aware In-context Construction
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by (Li and Pan, 2016; Zou et al., 2023). The optimization objective definition:
$\mathcal{T}^{\star}\!\!=\!\arg\!\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(\mathcal{G}),$ (1)
$\mathcal{H}^{\mathcal{T}}(\mathcal{G})=\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\!\mathcal{H}^{\mathcal{T}}(\mathcal{G},\phi)=-\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\frac{g_{\phi}}{\operatorname{vol}(\mathcal{G})}\log\!\frac{\operatorname{vol}\left(\phi\right)}{\operatorname{vol}\left(\phi^{+}\right)},$ (2)
@@
Theorem 4.1 (Topological Information Capturing Properties of Encoding Tree).
@@
Theorem 4.2 (Implicit Global Constraints in Low-Level Communities).
@@
Drawing on Theorems 4.1 and  4.2, we infer that low-level communities in the encoding tree provide a more faithful representation of the graph’s topology, as they simultaneously preserve local connectivity and global structural constraints. Accordingly, for each leaf node $\alpha$, we assign its associated low-level community $\mathcal{C}^{\ell}_{\alpha^{+}}$ as a topological context to enhance the accuracy of downstream inference by the LLMs. The proof process of Theorems 4.1 and  4.2 is provided in Appendix F.1-F.2.
@@ 4.2 Tree-prompted LLM Inference
Motivation. Based on the structural encoding tree $\mathcal{T}$, we integrate node language descriptions to enable comprehensive in-context learning with LLMs, improving their understanding of both topology and text without fine-tuning. The resulting inferences are then used to refine the encoding tree, strengthening homophily in low-level communities and providing a solid basis for subsequent GSL.
@@
$t^{\star}_{\alpha}=\operatorname{Con}\Big(\{t_{\alpha}\}\cup\{t_{j}:j\in\text{Top}_{k}(\mathbf{w}_{\alpha}),\ w_{\alpha j}\geq\epsilon\}\Big),$ (3)
$w_{\alpha\beta}=\frac{\exp\left(\operatorname{sim}(x_{\alpha},x_{\beta})\right)}{\sum\limits_{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}\exp\left(\operatorname{sim}(x_{\alpha},x_{\gamma})\right)},$ (4)
where $\operatorname{Con}$ denotes text concatenation, $\mathbf{w}_{\alpha}=\{w_{\alpha\beta}\}_{\beta\in\mathcal{C}^{\ell}_{\alpha^{+}}}$ and $\operatorname{sim}(x_{i},x_{j})$ represents the cosine similarity.
Community of Thought. Building on the enhanced leaf nodes, we introduce a Community of Thought (CoT) prompting mechanism, where each node is presented with community-aware descriptions and selected neighbors. This allows the LLM to jointly exploit topological and semantic information for more reliable inference. Concretely, CoT enables the LLM (i) to infer a label distribution for each node and (ii) to produce high-quality embeddings in its semantic space, which are then used to refine the structural encoding tree. In this way, CoT serves as an efficient in-context interface to the structure-aware encoding tree. Formally, given a target leaf node $\alpha$, the LLM inference with CoT can be written as:
$\mathbf{z}_{\alpha}=\Psi\bigl(\mathrm{LLM\mbox{-}CoT}(\mathcal{D}_{\text{task}},\mathcal{D}_{\text{topo}},\mathcal{D}_{\text{sema}},t^{\star}_{\alpha})\bigr),$ (5)
$y^{\mathrm{cls}}_{\alpha}\sim\mathrm{Categorical}\big(\operatorname{softmax}(\mathbf{z}_{\alpha})\big),$ (6)
where $\Psi(\cdot)$ is the extraction function that maps the LLM’s textual answer $\text{ANS}_{\alpha}$ into a numerical logit vector $\mathbf{z}_{\alpha}\in\mathbb{R}^{N_{\text{class}}}$, $y^{cls}_{\alpha}\in\Delta^{N_{\text{class}}-1}$ denotes the soft label probability simplex for node $\alpha$. Here, $\mathcal{D}_{\text{task}}$ specifies the task, while $\mathcal{D}_{\text{topo}}$ and $\mathcal{D}_{\text{sema}}$ encode topological and semantic context. An illustration of prompt construction is provided in Fig. 2.
@@
Based on Theorem 3, the similarity of soft labels $y^{\mathrm{cls}}$ approximates that of ground-truth labels $y$ within a controllable bias $\varepsilon^{y}(\alpha,\beta)$. This offers a solid theoretical foundation for the subsequent Leaf Dependency Allocation and Semantic Similarity Sampling.
@@
$\mathcal{P}^{\star}=\arg\max_{\mathcal{P}}\Bigg\{\sum_{\mathcal{C}\in\mathcal{P}}\max_{c}\sum_{v\in\mathcal{C}}y^{cls}_{v,c}+\beta\,\sum_{\mathcal{C}\in\mathcal{P}}\mathrm{Sil}(\mathcal{C};\mathbf{Y}^{cls})\Bigg\},$ (7)
$\mathcal{T}^{\star}=\mathrm{UpdateTree}(\mathcal{T},\mathcal{P}^{\star}).$ (8)